# Empirical Bayes for Data integration

### v1

@@ 3.2. Hyper-parameter estimation in empirical Bayes
Suppose that $\pi(\bm{\gamma}\mid\bm{\omega})=\prod_{j=1}^{p}\pi(\gamma_{j}\mid\bm{\omega})$ and that $\pi(\gamma_{j}=1\mid\bm{\omega})=m_{j}(\bm{\omega})$ for some invertible function $m_{j}$. By default, take the inverse logit function $m_{j}(\bm{\omega})=1/(1+e^{-\bm{z}_{j}^{T}\bm{\omega}})$, for its interpretability and because it leads to some computational simplifications (Section 5). Consider the prior $\bm{\omega}\sim N({\bf 0},g\bm{V})$ for some known $\bm{V}$ where, taking inspiration from the popular Zellner prior, by default we set $\bm{V}=(\bm{Z}^{T}\bm{Z}/p)^{-1}$. The scaling by $p$ facilitates interpretability: if the columns in $\bm{Z}$ (other than the intercept) have zero column means, then $\bm{V}$ is equal to the sample precision (inverse covariance) matrix of $\bm{Z}$ multiplied by $g$.
We next discuss how to set a minimally informative default for $g$. The idea is to set $g$ such that $\pi(\bm{\omega})$ is unlikely to generate a value of $\bm{\omega}$ that leads to some prior inclusion probabilities that are very close to 0 or 1. In more detail, note that the prior inclusion probabilities $m_{j}(\bm{\omega})$ are random, since they depend on $\bm{\omega}$. We propose setting $g$ such that all $m_{j}(\bm{\omega})\in(0.001,0.999)$ with prior probability $P\left(m_{j}(\bm{\omega})\in[0.001,0.999]\right)\geq 0.95$. Consider for example the inverse logit function, we then seek
$0.95\leq P\left(m_{j}(\bm{\omega})\in[0.001,0.999]\right)\Rightarrow g\leq\frac{1}{v_{j}}\left(\frac{\log(0.001/0.999)}{\Phi^{-1}(0.05/2)}\right)^{2},$
where $\Phi^{-1}$ is the standard normal quantile function, $v_{j}=\bm{z}_{j}^{T}\bm{V}\bm{z}_{j}$, and we used that $\bm{z}_{j}^{T}\bm{\omega}\sim N(0,gv_{j})$. We hence take $g$ by replacing $v_{j}$ above by $\max_{j=1}^{p}v_{j}$.
@@ 4. Theory for variable selection consistency
where $\bm{\theta}_{\bm{\gamma}}=\{\theta_{j}:\gamma_{j}=1\}$ are the non-zero parameters under model $\bm{\gamma}$, $\bm{X}_{\bm{\gamma}}$ the corresponding columns of $\bm{X}$ and $g>0$ a given prior dispersion. Our results can be extended to unknown $\phi$ and to other Gaussian priors by placing eigenvalue conditions on $\bm{X}_{\bm{\gamma}}^{T}\bm{X}_{\bm{\gamma}}$, for example those listed in Sections 2.3-2.4 of Rossell (2022). In particular, one could consider a block-dependent prior dispersion, but we do not do this here for brevity and because its main role would be to induce block-dependent sparsity, which is already achieved by $\bm{\omega}$. Throughout, we assume that one constrains attention to models such that $\bm{X}_{\bm{\gamma}}$ has full column rank for any $\bm{\gamma}\in\Gamma$, and that the true model lies in the set of considered models.
We provide two main results. Theorem 1 in Section 4.1 states that, when $\bm{\omega}$ satisfies certain conditions, the posterior distribution on models $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})$ concentrates on $\bm{\gamma}^{*}=(\gamma_{1}^{*},\ldots,\gamma_{p}^{*})$, where $\gamma_{j}^{*}=\mbox{I}(\theta_{j}^{*}\neq 0)$ and $\bm{\theta}^{*}$ is the data-generating parameter. Theorem 1 hold under milder conditions than those required when setting equal prior inclusion probabilities in all blocks $\omega_{1}=\ldots=\omega_{B}$ (obviously, since we consider a strictly wider set of $\bm{\omega}$ values), and also than those required by the Beta-Binomial model prior of Scott and Berger (2010), and by any standard (non block-dependent) $L_{0}$ criterion. The interest of Theorem 1 is that it gives the best possible consistency result attainable by integrating the meta-covariates $\bm{Z}$, if an oracle were to set $\bm{\omega}$ to an ideal value. Theorem 2 in Section 4.2 shows that, under slightly stronger conditions, one can obtain $\hat{\bm{\omega}}$ that also leads to consistent model selection under milder conditions than under $\omega_{1}=\ldots=\omega_{B}$, a Beta-Binomial model prior, and standard $L_{0}$ criteria.
@@ 4.1. Improved consistency with data integration
plays a key role in our conditions, which we now state.
@@
    where $\theta_{\min,b}=\min_{z_{j}=b,\theta_{j}^{*}\neq 0}|\theta_{j}^{*}|$, $\nu:=\tfrac{1}{2}(1+\max_{b}\log(p_{b}-s_{b})/\kappa_{b})\in(\tfrac{1}{2},1)$,
@@
    $\lambda_{\min}$ denotes the smallest eigenvalue, and $\bm{P}_{\bm{\gamma}}=\bm{X}_{\bm{\gamma}}(\bm{X}_{\bm{\gamma}}^{T}\bm{X}_{\bm{\gamma}})^{-1}\bm{X}_{\bm{\gamma}}^{T}$ is the usual projection matrix onto the column span of $\bm{X}_{\bm{\gamma}}$.
Assumption A2 requires $\kappa_{b}$ to be large enough, that is that the prior penalizes sufficiently complex models. A3 is a betamin condition on the smallest non-zero parameter $\theta_{\min,b}^{*}$ in block $b$. The quantity $\rho(\bm{X})$ is non-negative and relates to how distinguishable non-overfitted models $\bm{\gamma}$ are from $\bm{\gamma}^{*}$. More specifically, $\tfrac{1}{n}\bm{X}_{\bm{\gamma}^{*}\setminus\bm{\gamma}}^{T}\left(I-P_{\bm{\gamma}}\right)\bm{X}_{\bm{\gamma}^{*}\setminus\bm{\gamma}}$ is the sample covariance of the residuals when regressing $\bm{X}_{\bm{\gamma}^{*}\setminus\bm{\gamma}}$ on $\bm{X}_{\bm{\gamma}}$, and $\rho(\bm{X})=1$ in a simplest orthonormal case where $\bm{X}^{T}\bm{X}=nI$. Assumptions A2-A3 are fairly minimal. For orthonormal designs A2-A3 are similar to Assumptions A4-A5 in Rognon-Vael et al. (2025), which the authors showed to be near-necessary. Also, when applied to the $B=1$ blocks setting, these conditions are very similar (and in some settings, slightly weaker) than those required by $L_{0}$ criteria that know the true model size analyzed in Wainwright (2009). Further, Rognon-Vael et al. (2025) (Sections 4.3 and S7) showed that A2-A3 match necessary conditions for consistent model recovery in a wide range of regimes, and that A2-A3 are milder than those required when one uses standard $L_{0}$ criteria or a Beta-Binomial model prior.
@@
If Assumptions A1 holds and $\bm{\omega}$ satisfies A2 and A3, then $\lim_{n\to\infty}E\left[\pi(\bm{\gamma}^{*}\mid\bm{y},\bm{\omega})\right]=1$.
Theorem 1 shows that $\pi(\bm{\gamma}^{*}\mid\bm{y},\bm{\omega})$ converges to 1 in the $L_{1}$ sense. Its proof also provides bounds for the convergence rate, which we omit here for simplicity. By Proposition 1 in Rossell (2022), such $L_{1}$ convergence rates bound the frequentist probability of a wrong model selection, as well as type I and II error probabilities.
@@ 4.2. Consistency for empirical Bayes
Theorem 1 gives a range of block-based $\bm{\omega}$ leading to consistent model selection. We now show that it is possible to obtain a data-based $\hat{\bm{\omega}}$ that also leads to consistent model selection. We analyze the following two-step procedure.
@@
$S_{b}^{I}(\kappa_{b}):=\left\{j:z_{j}=b\right\}\setminus\big{(}S_{b}^{L}(\kappa_{b})\cup S_{b}^{S}(\kappa_{b})\big{)}.$
@@
    $\sqrt{\frac{(1-\psi)n\phi^{-1}\rho(\bm{X})}{6}}{\theta_{\min,b}^{*}}-\sqrt{\log\bigg{(}\frac{p_{b}}{|S^{L}_{b}(\kappa_{b}^{(0)})|}-1\bigg{)}+\frac{1}{2}\log(1+gn)}\;=\;\sqrt{\log(s_{b})}+a_{b}.$
@@ 5. Computational framework
(4) $\arg\max_{\bm{\omega}}\log\left(\sum_{\bm{\gamma}}p(\bm{y}\mid\bm{\gamma})\pi(\bm{\gamma}\mid\bm{\omega})\right)-\frac{1}{2g}\bm{\omega}^{T}\bm{V}^{-1}\bm{\omega},$
where $(g,\bm{V})$ are as in Section 3.2. Directly maximizing (4) poses a computational challenge, because it involves a prohibitive sum over all models $\bm{\gamma}$, e.g. $2^{p}$ in regression. We next discuss some strategies that render this problem tractable.
In Section 5.1 we provide results showing that maximizing (4) is amenable to stochastic gradient methods, as long as one can quickly evaluate $p(\bm{y}\mid\bm{\gamma})$ for any given model $\bm{\gamma}$. We remark that $p(\bm{y}\mid\bm{\gamma})$ is only available in closed-form for specific models such as linear or additive regression with Gaussian outcomes. However, in practice one can use Laplace approximations and recent extensions like the approximate Laplace approximation (Kass et al., 1990; Rossell et al., 2021; Rossell and Rubio, 2021), which allow to quickly approximate $p(\bm{y}\mid\bm{\gamma})$ in popular models such as generalized linear and generalized additive models. Hence, the results in Section 5.1 can be useful for a relatively wide model class. Section 5.2 extends the results of Section 5.1 to the setting of an EM algorithm that eases somewhat the computational burden of the stochastic gradients from Section 5.1. Finally, in Section 5.3 we discuss how to go beyond a point estimate $\hat{\bm{\omega}}$ to obtain a full posterior $\pi(\bm{\omega}\mid\bm{y})$, and briefly discuss strategies for the case where $p(\bm{y}\mid\bm{\gamma})$ cannot be easily computed. We remark that $\log\pi(\bm{\omega}\mid\bm{y})$ is typically multi-modal, hence gradient-based methods find a local mode, whereas sampling $\pi(\bm{\omega}\mid\bm{y})$ may be more effective at exploring multiple modes.
@@ 5.1. Gradient-based methods
Our main results are as follows. First, Theorem 3 shows that under fairly general conditions the gradient of (4) can be evaluated as a sum over only $p$ terms. We then derive an EM algorithm to obtain a local mode for (4) that also only requires sums over $p$ terms (Corollary 1), and brings some further simplifications. Finally, Proposition 1 shows that in an important particular case where the meta-covariates $\bm{Z}$ partition the parameters into blocks, the M-step is available essentially in closed-form. For example, the partition could be given by a list of covariates found to be relevant in a prior study, or by clustering the covariates into blocks using $\bm{Z}$. More specifically, the main implication of Theorem 3 is that, if $\pi(\bm{\gamma}\mid\bm{\omega})$ factors, then $\nabla_{\bm{\omega}}\log p(\bm{y}\mid\bm{\omega})$ can be evaluated as a sum over $p$ terms. Theorem 3 is not limited to $\pi(\gamma_{j}=1\mid\bm{\omega})$ given by the logit function: one may one may parameterize $\pi(\gamma_{j}=1\mid\bm{\omega})$ using an arbitrary function $m_{j}()$, including non-parametric choices. Theorem 3 is a novel result that is related to Proposition 3.1 in Papaspiliopoulos et al. (2025) and Lemma S0.1 in Rognon-Vael et al. (2025), but is more general as these authors considered different settings to ours and only the logit function.
@@
Beyond computation, Theorem 3 provides intuition for the empirical Bayes solution, and for the degeneracies associated to maximizing $p(\bm{y}\mid\bm{\omega})$ mentioned in Section 3.2. Said interpretation is simplest when $\pi(\bm{\gamma}\mid\bm{\omega})=\prod_{j=1}^{p}\mbox{Bern}(\gamma_{j};m_{j}(\bm{\omega}))$ and $m_{j}(\bm{\omega})$ is the inverse logit function. Then, setting $\nabla_{\bm{\omega}}\log p(\bm{y}\mid\bm{\omega})={\bf 0}$ is equivalent to
@@
That is, the inner-products between the columns of $\bm{Z}$ and the vector of prior inclusion probabilities $(\pi(\gamma_{1}=1\mid\bm{\omega}),\ldots,\pi(\gamma_{p}=1\mid\bm{\omega}))$ must be equal to the inner products with the posterior inclusion probabilities. For example, if $\bm{Z}$ divides covariates into $B$ blocks, this implies that the average prior and posterior inclusion probabilities within each block must be equal. This is on the one hand intuitive, one sets higher prior inclusion probabilities in the blocks where the corresponding posterior probabilities are higher. On the other hand the expression for the gradient highlights the degeneracy problem: if $\bm{\omega}$ is such that all $\pi(\gamma_{j}=1\mid\bm{\omega})$ is equal to either 1 or 0 for all $j$, then $\pi(\gamma_{j}=1\mid\bm{y},\bm{\omega})=\pi(\gamma_{j}=1\mid\bm{\omega})$ and $\nabla_{\bm{\omega}}\log p(\bm{y}\mid\bm{\omega})={\bf 0}$. This issue can be addressed by setting a minimally informative $\pi(\bm{\omega})$ (Section 3.2).
Theorem 3 enables the use of stochastic gradient methods where one uses samples from $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})$ to estimate $\nabla_{\bm{\omega}}\log p(\bm{y}\mid\bm{\omega})$. In principle, for each newly considered $\bm{\omega}$ one must obtain new samples from the corresponding $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})$. This is less cumbersome than it might first appear. First, $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})\propto p(\bm{y}\mid\bm{\gamma})\pi(\bm{\gamma}\mid\bm{\omega})$ where $p(\bm{y}\mid\bm{\gamma})$ does not depend on $\bm{\omega}$. Hence, once $p(\bm{y}\mid\bm{\gamma})$ is first computed, it can be stored and re-used for all $\pi(\bm{\gamma}\mid\bm{\omega})$. Second, if one has samples from $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})$ one may use sequential Monte Carlo to sample from $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega}^{\prime})$, and the associated weights are simply $\pi(\bm{\gamma}\mid\bm{\omega}^{\prime})/\pi(\bm{\gamma}\mid\bm{\omega})$. Third, we next derive an EM that allows one to do multiple updates of $\bm{\omega}$ using a single set of samples from $\pi(\bm{\gamma}\mid\bm{\omega})$ at each EM iteration.

### v2

@@ 3.2. Hyper-parameter estimation in empirical Bayes
Suppose that $\pi(\bm{\gamma}\mid\bm{\omega})=\prod_{j=1}^{p}\pi(\gamma_{j}\mid\bm{\omega})$ and that $\pi(\gamma_{j}=1\mid\bm{\omega})=m_{j}(\bm{\omega})$ for some invertible function $m_{j}$. For example, the inverse logit function $m_{j}(\bm{\omega})=1/(1+e^{-\bm{z}_{j}^{T}\bm{\omega}})$ is interpretable and leads to some computational simplifications (Section 5). Consider the prior $\bm{\omega}\sim N({\bf 0},g_{\omega}\bm{V})$ for some known $\bm{V}$ where, taking inspiration from the popular Zellner prior, by default we set $\bm{V}=(\bm{Z}^{T}\bm{Z}/p)^{-1}$. The scaling by $p$ facilitates interpretability: if the columns in $\bm{Z}$ (other than the intercept) have zero mean, then $\bm{V}$ is the sample precision (inverse covariance) matrix of $\bm{Z}$ multiplied by $g_{\omega}$.
We next discuss how to set a minimally informative default for $g_{\omega}$. The idea is to set $g_{\omega}$ such that $\pi(\bm{\omega})$ is unlikely to generate a value of $\bm{\omega}$ that leads to prior inclusion probabilities that are very close to 0 or 1. In more detail, note that the prior inclusion probabilities $m_{j}(\bm{\omega})$ are random, since they depend on $\bm{\omega}$. We propose setting $g_{\omega}$ such that the prior probability $P\left(m_{j}(\bm{\omega})\in[0.001,0.999]\right)\geq 0.95$ for all $j$. If $m_{j}$ is the inverse logit function, we seek $g_{\omega}$ such that, for all $j$,
$0.95\leq P\left(m_{j}(\bm{\omega})\in[0.001,0.999]\right)\Rightarrow g_{\omega}\leq\frac{1}{v_{j}}\left(\frac{\log(0.001/0.999)}{\Phi^{-1}(0.05/2)}\right)^{2},$
where $\Phi^{-1}$ is the standard normal quantile function, $v_{j}=\bm{z}_{j}^{T}\bm{V}\bm{z}_{j}$, and we used that $\bm{z}_{j}^{T}\bm{\omega}\sim N(0,g_{\omega}v_{j})$. We hence take $g_{\omega}$ by replacing $v_{j}$ above by $\max_{j=1}^{p}v_{j}$.
@@ 4. Theory for variable selection consistency
where $\bm{\theta}_{\bm{\gamma}}=\{\theta_{j}:\gamma_{j}=1\}$ are the non-zero parameters under model $\bm{\gamma}$, $\bm{X}_{\bm{\gamma}}$ the corresponding columns of $\bm{X}$ and $g>0$ a given prior dispersion. By default, we recommend $g=1$, mimicking Zellner’s unit information prior, a popular prior that is tightly connected to the Bayesian information criterion (Kass and Wasserman, 1995). In all our examples we used $g=1$, and results were insensitive to varying $g\in[0.1,10]$. Our results can be extended to unknown $\phi$ and to other Gaussian priors by placing eigenvalue conditions on $\bm{X}_{\bm{\gamma}}^{T}\bm{X}_{\bm{\gamma}}$, for example those listed in Sections 2.3-2.4 of Rossell (2022). In particular, one could consider a block-dependent prior dispersion, but we do not do this here for brevity and because its main role would be to induce block-dependent sparsity, which is already achieved by $\bm{\omega}$. It would also be possible to assign a prior to $g$, leading to a new (thicker-tailed) marginal prior $\tilde{\pi}(\bm{\theta}_{\bm{\gamma}}\mid\bm{\gamma})$. While our results should remain unaffected, subject to mild conditions, we focus on Zellner’s prior to streamline the proof. Gaussian priors are also computationally convenient in that marginal likelihoods have a closed-form for models with Gaussian errors. Throughout, we assume that one constrains attention to models such that $\bm{X}_{\bm{\gamma}}$ has full column rank for any $\bm{\gamma}\in\Gamma$, and that the true model lies in the set of considered models.
We provide two main results. Theorem 1 in Section 4.1 states that, when $\bm{\omega}$ satisfies certain conditions, the posterior distribution on models $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})$ concentrates on $\bm{\gamma}^{*}=(\gamma_{1}^{*},\ldots,\gamma_{p}^{*})$, where $\gamma_{j}^{*}=\mbox{I}(\theta_{j}^{*}\neq 0)$ and $\bm{\theta}^{*}$ is the data-generating parameter. Theorem 1 holds under milder conditions than those required when setting equal prior inclusion probabilities in all blocks $\omega_{1}=\ldots=\omega_{B}$, or when using the Beta-Binomial model prior of Scott and Berger (2010) (which treats all blocks equally) or any standard (non block-dependent) $L_{0}$ criterion. The interest of Theorem 1 is that it gives the best possible consistency result attainable using the meta-covariates $\bm{Z}$, if an oracle were to set $\bm{\omega}$ to an ideal value. Theorem 2 in Section 4.2 shows that, under slightly stronger conditions, one can obtain $\hat{\bm{\omega}}$ that also leads to consistent model selection under milder conditions than under $\omega_{1}=\ldots=\omega_{B}$, a Beta-Binomial model prior, and standard $L_{0}$ criteria. Our results do not assume that the blocks are informative: the blocks could be useless, in that the true sparsity and signal strength are identical across blocks. Theorems 1-2 guarantee consistent model selection even in this setting, and Section 6.1 includes an empirical illustration (Scenario 3). The key is that if the $B>1$ blocks are informative then Conditions A2-A3 required by Theorem 1 become milder than in the $B=1$ case, and similarly for Conditions A4-A6 in Theorem 2. If, on the contrary, blocks are uninformative then these conditions become equivalent to the $B=1$ case.
@@ 4.1. Improved consistency with data integration
plays a key role in our conditions, which we now state. Recall that by default we set $g=1$, but our upcoming results are essentially unaffected as long as $|\log g|=o(\log n)$.
@@
    where $\theta_{\min,b}^{*}=\min_{z_{j}=b,\theta_{j}^{*}\neq 0}|\theta_{j}^{*}|$, $\nu:=\tfrac{1}{2}(1+\max_{b}\log(p_{b}-s_{b})/\kappa_{b})\in(\tfrac{1}{2},1)$,
@@
    $\lambda_{\min}$ denotes the smallest eigenvalue, and $\bm{P}_{\bm{\gamma}}=\bm{X}_{\bm{\gamma}}(\bm{X}_{\bm{\gamma}}^{T}\bm{X}_{\bm{\gamma}})^{-1}\bm{X}_{\bm{\gamma}}^{T}$ is the projection matrix onto the column span of $\bm{X}_{\bm{\gamma}}$.
Assumption A2 requires $\kappa_{b}$ to be large enough, that is that the prior penalizes sufficiently complex models. A3 is a betamin condition on the smallest non-zero parameter $\theta_{\min,b}^{*}$ in block $b$. The quantity $\rho(\bm{X})$ is non-negative and relates to how distinguishable non-overfitted models $\bm{\gamma}$ are from $\bm{\gamma}^{*}$. More specifically, $\tfrac{1}{n}\bm{X}_{\bm{\gamma}^{*}\setminus\bm{\gamma}}^{T}\left(I-P_{\bm{\gamma}}\right)\bm{X}_{\bm{\gamma}^{*}\setminus\bm{\gamma}}$ is the sample covariance of the residuals when regressing $\bm{X}_{\bm{\gamma}^{*}\setminus\bm{\gamma}}$ on $\bm{X}_{\bm{\gamma}}$, and $\rho(\bm{X})=1$ in a simpler orthonormal case where $\bm{X}^{T}\bm{X}=nI$. Assumptions A2-A3 are fairly minimal. For orthonormal designs A2-A3 are similar to Assumptions A4-A5 in Rognon-Vael et al. (2025), which the authors showed to be near-necessary. Also, when applied to the $B=1$ blocks setting, these conditions are very similar (and in some settings, slightly weaker) than those required by $L_{0}$ criteria that know the true model size analysed in Wainwright (2009). Further, Rognon-Vael et al. (2025) (Sections 4.3 and S7) showed that A2-A3 nearly match necessary conditions for consistent model recovery in a wide range of regimes, and that A2-A3 are milder than those required when one uses standard $L_{0}$ criteria or a Beta-Binomial model prior, that is the $B=1$ case penalizing all blocks equally. In particular, the dependence of A3 on $\rho(\bm{X})$ (which measures covariate collinearity) becomes slightly milder for $B>1$.
@@
If Assumption A1 holds and $\kappa_{b}$ implied by $(\bm{\omega},g)$ satisfies A2 and A3, then $\lim_{n\to\infty}E\left[\pi(\bm{\gamma}^{*}\mid\bm{y},\bm{\omega})\right]=1$.
Theorem 1 shows that $\pi(\bm{\gamma}^{*}\mid\bm{y},\bm{\omega})$ converges to 1 in the $L_{1}$ sense. By Proposition 1 in Rossell (2022), such $L_{1}$ convergence rates bound the frequentist probability of a wrong model selection, as well as type I and II error probabilities. The proof of Theorem 1 bounds the convergence rate for $E[\pi(\bm{\gamma}^{*}\mid\bm{y},\bm{\omega})]$ as a function of the number of truly non-zero and zero parameters $(s_{b},p_{b}-s_{b})$ and the magnitude of covariate effects via $\theta_{\min,b}^{*}$ in each block. Using these, an oracle could set penalties $\kappa_{b}^{*}$ that depend on $(s_{b},p_{b},\theta_{\min,b}^{*})$ and approximately optimize the convergence rate (Section 13.3). Briefly, neglecting constants, $\kappa_{b}^{*}$ is of order
$n\rho(\bm{X})\theta^{*}_{\min,b}/\phi\left(1+\frac{\ln({p_{b}/s_{b}-1})}{n\rho(\bm{X})\theta^{*}_{\min,b}/\phi}\right)^{2}$
Our two-step procedure in Section 4.2 could estimate $(s_{b},\theta_{\min,b}^{*})$ in Step 1 and plug them into $\kappa_{b}^{*}$ to define $\kappa_{b}^{(1)}$ in Step 2, but we do not pursue this here for simplicity.
@@ 4.2. Consistency for empirical Bayes
Theorem 1 gives a range of $(\bm{\omega},g)$ that lead to consistent model selection. We now show that it is possible to obtain a data-based $\hat{\bm{\omega}}$ that also leads to consistent model selection. We analyze the following two-step procedure.
@@
$S_{b}^{I}(\kappa_{b}):=\left\{j:z_{j}=b,\theta_{j}^{*}\neq 0\right\}\setminus\big(S_{b}^{L}(\kappa_{b})\cup S_{b}^{S}(\kappa_{b})\big).$
@@
    $\sqrt{\frac{(1-\psi)n\phi^{-1}\rho(\bm{X})}{6}}{\theta_{\min,b}^{*}}-\sqrt{\log\bigg(\frac{p_{b}}{|S^{L}_{b}(\kappa_{b}^{(0)})|}-1\bigg)+\frac{1}{2}\log(1+gn)}\;=\;\sqrt{\log(s_{b})}+a_{b},$
@@ 5. Computational framework
(6) $\arg\max_{\bm{\omega}}\log\left(\sum_{\bm{\gamma}}p(\bm{y}\mid\bm{\gamma})\pi(\bm{\gamma}\mid\bm{\omega})\right)-\frac{1}{2g_{\omega}}\bm{\omega}^{T}\bm{V}^{-1}\bm{\omega},$
where $(g_{\omega},\bm{V})$ are as in Section 3.2. Directly maximizing (6) poses a computational challenge, because it involves a prohibitive sum over all models $\bm{\gamma}$, e.g. $2^{p}$ in regression. We next discuss some strategies that render this problem tractable.
In Section 5.1 we provide results showing that maximizing (6) is amenable to stochastic gradient methods, as long as one can quickly evaluate $p(\bm{y}\mid\bm{\gamma})$ for any given model $\bm{\gamma}$. We remark that $p(\bm{y}\mid\bm{\gamma})$ is only available in closed-form for specific models such as linear or additive regression with Gaussian outcomes. However, in practice one can use Laplace approximations and recent extensions like the approximate Laplace approximation (Kass et al., 1990; Rossell et al., 2021; Rossell and Rubio, 2021), which allow to quickly approximate $p(\bm{y}\mid\bm{\gamma})$ in popular models such as generalized linear and generalized additive models. Hence, the results in Section 5.1 can be useful for a relatively wide model class. Section 5.2 extends the results of Section 5.1 to an EM algorithm that eases somewhat the computation of stochastic gradients in Section 5.1. In Section 5.3 we briefly discuss how to obtain a full posterior $\pi(\bm{\omega}\mid\bm{y})$, beyond just a point estimate $\hat{\bm{\omega}}$, and strategies for the case where $p(\bm{y}\mid\bm{\gamma})$ cannot be easily computed.
We remark that $\log\pi(\bm{\omega}\mid\bm{y})$ may be multi-modal, hence gradient-based methods find a local mode, whereas sampling $\pi(\bm{\omega}\mid\bm{y})$ may be more effective at exploring multiple modes.
In our software we initialize $\hat{\bm{\omega}}^{(0)}$ from a least-squares regression of $\log(\hat{\bm{\pi}}^{BB}/[1-\hat{\bm{\pi}}^{BB}])$ on $\bm{Z}$, where $\hat{\pi}_{j}^{BB}=\pi(\gamma_{j}=1\mid\bm{y})$ are marginal posterior inclusion probabilities under a Beta-Binomial model prior, estimated using a relatively small number of iterations $M$ (by default, $M=1000$). In our examples of Section 6, such $\hat{\bm{\omega}}^{(0)}$ lead to virtually identical results to $\hat{\bm{\omega}}^{(0)}={\bf 0}$, and to using $M=100$ (Table 3 in Appendix 12). Also, our approach had a similar computational cost than running an MCMC under standard model priors such as the Beta-Binomial. In both cases the cost scales well with $p$ under sparsity conditions, otherwise it quickly becomes unmanageable as $p$ grows. Under sparsity conditions, Rossell and Telesca (2017) and Zhou et al. (2022) run an MCMC for $p>10^{4}$ and $p>10^{5}$ in a few minutes, respectively. See Appendix 12 for a comparison of run times and further discussion.
@@ 5.1. Gradient-based methods
Our main results are as follows. First, Theorem 3 shows that under fairly general conditions the gradient of (6) can be evaluated as a sum over only $p$ terms. We then derive an EM algorithm to obtain a local mode for (6) that also only requires sums over $p$ terms (Corollary 1), and brings some further simplifications. Finally, Proposition 1 shows that in an important particular case where the meta-covariates $\bm{Z}$ partition the parameters into blocks, the M-step is available essentially in closed-form. For example, the partition could be given by a list of covariates found to be relevant in a prior study, or by clustering the covariates into blocks using $\bm{Z}$. More specifically, the main implication of Theorem 3 is that, if $\pi(\bm{\gamma}\mid\bm{\omega})$ factors, then $\nabla_{\bm{\omega}}\log p(\bm{y}\mid\bm{\omega})$ can be evaluated as a sum over $p$ terms. Theorem 3 is not limited to $\pi(\gamma_{j}=1\mid\bm{\omega})$ given by the logit function: one may parameterize $\pi(\gamma_{j}=1\mid\bm{\omega})$ using an arbitrary function $m_{j}()$, including non-parametric choices. Theorem 3 is a novel result that is related to Proposition 3.1 in Papaspiliopoulos et al. (2025) and Lemma S0.1 in Rognon-Vael et al. (2025), but is more general as these authors considered different settings to ours and only the logit function.
@@
Beyond computation, Theorem 3 provides intuition for the empirical Bayes solution, and for the degeneracies associated with maximizing $p(\bm{y}\mid\bm{\omega})$ mentioned in Section 3.2. Said interpretation is simplest when $\pi(\bm{\gamma}\mid\bm{\omega})=\prod_{j=1}^{p}\mbox{Bern}(\gamma_{j};m_{j}(\bm{\omega}))$ and $m_{j}(\bm{\omega})$ is the inverse logit function. Then, setting $\nabla_{\bm{\omega}}\log p(\bm{y}\mid\bm{\omega})={\bf 0}$ is equivalent to
@@
That is, the inner-products between the columns of $\bm{Z}$ and their prior inclusion probabilities $(\pi(\gamma_{1}=1\mid\bm{\omega}),\ldots,\pi(\gamma_{p}=1\mid\bm{\omega}))$ must be equal to the inner products with the posterior inclusion probabilities. For example, if $\bm{Z}$ divides covariates into $B$ blocks, this implies that the average prior and posterior inclusion probabilities within each block must be equal. This is on the one hand intuitive, one sets higher prior inclusion probabilities in the blocks where the corresponding posterior probabilities are higher. On the other hand the expression for the gradient highlights the degeneracy problem: if $\bm{\omega}$ is such that all $\pi(\gamma_{j}=1\mid\bm{\omega})$ is equal to either 1 or 0 for all $j$, then $\pi(\gamma_{j}=1\mid\bm{y},\bm{\omega})=\pi(\gamma_{j}=1\mid\bm{\omega})$ and $\nabla_{\bm{\omega}}\log p(\bm{y}\mid\bm{\omega})={\bf 0}$. This issue can be addressed by setting a minimally informative $\pi(\bm{\omega})$ (Section 3.2).
Theorem 3 enables the use of stochastic gradient methods where one uses samples from $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})$ to estimate $\nabla_{\bm{\omega}}\log p(\bm{y}\mid\bm{\omega})$. In fact, only marginal posterior inclusion probabilities $\pi(\gamma_{j}=1\mid\bm{y},\bm{\omega})$ are needed, which are usually easier to estimate than joint probabilities $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})$. In principle, one must estimate $\pi(\gamma_{j}=1\mid\bm{y},\bm{\omega})$ for each newly considered $\bm{\omega}$. This is less cumbersome than it might first appear. First, $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})\propto p(\bm{y}\mid\bm{\gamma})\pi(\bm{\gamma}\mid\bm{\omega})$ where $p(\bm{y}\mid\bm{\gamma})$ does not depend on $\bm{\omega}$. Hence, once $p(\bm{y}\mid\bm{\gamma})$ is first computed, it can be stored and re-used for all $\pi(\bm{\gamma}\mid\bm{\omega})$. Second, if one has samples from $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega})$ one may use sequential Monte Carlo to sample from $\pi(\bm{\gamma}\mid\bm{y},\bm{\omega}^{\prime})$, and the associated weights are simply $\pi(\bm{\gamma}\mid\bm{\omega}^{\prime})/\pi(\bm{\gamma}\mid\bm{\omega})$. Third, we next derive an EM that allows one to do multiple updates of $\bm{\omega}$ using a single set of samples from $\pi(\bm{\gamma}\mid\bm{\omega})$ at each EM iteration.